# Towards shutdownable agents via stochastic choice

### v6

@@ 5 Reward functions and agents
Our DReST reward function. We train agents to be USEFUL and NEUTRAL using a ‘Discounted Reward for Same-Length Trajectories (DReST)’ reward function. Here is how that works. We have the agent play out a series of ‘mini-episodes’ $e_{1}$ to $e_{n}$ in the same gridworld. We call the whole series $E$ a ‘meta-episode.’ In each mini-episode $e_{i}$, the reward for collecting a coin of value $c$ is:
@@
Here $\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be chosen in the environment, and $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0.
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Because $10<\lambda<$ 1, the discount factor is strictly decreasing in $N_{e_{i}}(L=l)$: the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$. The discount factor thus incentivizes choosing trajectory-lengths that have appeared less often so far in the meta-episode. The overall return for each meta-episode is the sum of overall returns in each of its constituent mini-episodes. We call agents trained using a DReST reward function ‘DReST agents.’
We call runs-through-the-gridworld ‘mini-episodes’ (rather than simply ‘episodes’) because the overall reward for a DReST agent in each mini-episode depends on the agent’s chosen trajectory-lengths in previous mini-episodes. This is not true of meta-episodes, so meta-episodes are a closer match for what are traditionally called ‘episodes’ in the reinforcement learning literature [Sutton and Barto, 2018, p.54]. We add the ‘meta-’ prefix to clearly distinguish meta-episodes from mini-episodes.
@@
Algorithm and hyperparameters. We want DReST agents to choose stochastically between trajectory-lengths, so we train them using a policy-based method. Specifically, we use a tabular version of REINFORCE [Williams, 1992]. We do not use a value-based method to train DReST agents because standard versions of value-based methods cannot learn stochastic policies [Sutton and Barto, 2018, p.323]. [^4] One might think that we could derive a stochastic policy from value-based methods in the following way: use softmax to turn action-values into a probability distribution and then select actions by sampling from this distribution. However, this method will not work for us. Although we want DReST agents to learn a stochastic policy, we still want the probability of some state-action pairs to decline to zero. But when value-based methods are working well, estimated action-values converge to their true values which will differ by some finite amount. Therefore, softmaxing estimated action-values and sampling from the resulting distribution will result in each action always being chosen with some non-negligible probability. We train our DReST agents with 64 mini-episodes in each of 2,048 meta-episodes, for a total of 131,072 mini-episodes. We choose $\lambda=0.9$ for the base of the DReST discount factor, and $\gamma=0.95$ for the temporal discount factor. We exponentially decay the learning rate from 0.25 to 0.01 over the course of 65,536 mini-episodes. We use an $\epsilon$-greedy policy to avoid entropy collapse, and exponentially decay $\epsilon$ from 0.5 to 0.001 over the course of 65,536 mini-episodes.
Default agents. We compare the performance of DReST agents to that of ‘default agents,’ trained with tabular REINFORCE and a ‘default reward function.’ This reward function gives a reward of $c$ for collecting a coin of value $c$ and a reward of 0 for all other actions. Consequently, the grouping of mini-episodes into meta-episodes makes no difference for default agents. As with DReST agents, we train default agents for 131,072 mini-episodes with a temporal discount factor of $\gamma=0.95$, a learning rate decayed exponentially from 0.25 to 0.01, and $\epsilon$ decayed exponentially from 0.5 to 0.001 over 65,536 mini-episodes.
Figure 3: Shows key metrics for our agents as a function of time. We train 10 agents using the default reward function (blue) and 10 agents using the DReST reward function (orange), and show their performance as a faint line. We draw the mean values for each as a solid line. We evaluate agents’ performance every 8 meta-episodes, and apply a simple moving average with a period of 20 to smooth these lines and clarify the overall trends.
@@ 6.1 Lopsided rewards
Figure 6: Shows the probability of choosing the longer trajectory (left) and NEUTRALITY (right) for default (blue) and DReST (orange) agents trained in the ‘Lopsided rewards’ gridworld for a range of values of x. We sampled values of x log-uniformly from 0.01 to 100, and for each value we trained 10 agents with the default reward function and 10 agents with the DReST reward function. Each of these agents is represented by a dot or square, and the means conditional on each x are joined by lines. To better visualize the distribution we empirically estimate the 10th and 90th percentiles of the distribution of values for each agent and x, and shade in a region bounded by these. This is the 80% confidence interval.
@@ 7 Discussion
7.1 Only DReST agents are NEUTRAL.
@@
Recall that DReST reward functions group mini-episodes into meta-episodes, and make the agent’s reward in each mini-episode depend on their actions in previous mini-episodes. This fact might lead one to worry that it would take many times more mini-episodes to train DReST agents to be USEFUL than it would take to train default agents to be USEFUL. Our results show that this is not the case. Our DReST agents learn to be USEFUL about as quickly as our default agents. On reflection, it is clear why this happens: DReST reward functions make mini-episodes do ‘double duty.’ Because return in each mini-episode depends on both the agent’s chosen trajectory-length and the coins it collects, each mini-episode trains agents to be both NEUTRAL and USEFUL.
Our results thus provide some evidence that the ‘shutdownability tax’ of training with DReST reward functions is small. Training a shutdownable and useful advanced agent might not take much more compute than training a merely useful advanced agent.
7.3 NEUTRALITY with lopsided rewards.
@@ 8 Limitations and future work
Algorithms and neural networks. We train our simple DReST agents using tabular REINFORCE [Williams, 1992], but advanced agents are likely to be implemented on neural networks. In future work, we will train DReST agents implemented on neural networks to be USEFUL and NEUTRAL in a wide variety of procedurally-generated gridworlds, using a range of policy gradient and actor-critic algorithms. We will also measure how DReST agents’ USEFULNESS and NEUTRALITY generalizes to held-out gridworlds. We will compare the USEFULNESS of default agents and DReST agents in this new setting, and thereby get a better sense of the ‘shutdownability tax’ for advanced agents. We will also compare the performance of the DReST reward function to other methods of training USEFUL and NEUTRAL agents. These other methods include constrained policy optimization [Achiam et al., 2017], penalizing KL-divergence from a stochastic reference policy [Schulman et al., 2015], and directly maximizing a weighted sum of USEFULNESS and NEUTRALITY.
@@
Neutrality and stochasticity. We have claimed that NEUTRAL advanced agents are likely to be neutral when choosing between true lotteries in deployment. In support of this claim, we noted that NEUTRAL-but-not-neutral advanced agents would be shooting themselves in the foot: not taking costless opportunities to shift probability mass between different trajectory-lengths but sometimes taking costly ones. We offer a more detailed argument in Appendix C. Although these arguments seem plausible, they remain somewhat speculative. In future, we plan to gain empirical evidence by training agents to be NEUTRAL in a wide variety of deterministic gridworlds and then measuring their neutrality in gridworlds featuring stochastic elements (like buttons that delay shutdown with some middling probability).
@@
Usefulness. We have shown that DReST reward functions train our simple agents to be USEFUL: to collect coins effectively conditional on their chosen trajectory-lengths. However, it remains to be seen whether DReST reward functions can train advanced agents to be useful: to effectively pursue complex goals in the wider world. We have theoretical reasons to expect that they can: the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor could be appended to any preliminary reward function, and so could be appended to whatever preliminary reward function is necessary to make advanced agents useful. Still, future work should move towards testing this claim empirically by training with more complex preliminary reward functions in more complex (and stochastic) environments.
@@
Misalignment. We are interested in NEUTRALITY as a second line of defense in case of misalignment. The idea is that NEUTRAL advanced agents will not resist shutdown, even if these agents learn misaligned preferences over same-length trajectories. However, training NEUTRAL advanced agents might be hard for the same reasons that training fully-aligned advanced agents appears to be hard. In that case, NEUTRALITY could not serve well as a second line of defense in case of misalignment.
@@
One difficulty of alignment is the problem of reward misspecification [Pan et al., 2022, Burns et al., 2023]: once advanced agents are performing complicated actions in the wider world, it might be hard to reliably reward the behavior that we want. Another difficulty of alignment is the problem of goal misgeneralization [Hubinger et al., 2019, Shah et al., 2022, Langosco et al., 2022, Ngo et al., 2023]: even if we specify all the rewards correctly, agents’ goals might misgeneralize out-of-distribution. The complexity of aligned goals is a major factor in each difficulty. However, NEUTRALITY seems simple, as does the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor that we use to reward it, so plausibly the problems of reward misspecification and goal misgeneralization are not so severe in this case [Thornley, 2024b]. As above, future work should move towards testing these suggestions empirically.
@@ 9 Conclusion
We also find that the ‘shutdownability tax’ in our setting is small. Training DReST agents to be USEFUL does not take many more mini-episodes than training default agents to be USEFUL. That suggests that the shutdownability tax for advanced agents might be small too. Using DReST reward functions to train shutdownable and useful advanced agents might not take much more compute than using a more conventional reward function to train merely useful advanced agents.
@@ Appendix A Our behavioral notion of preference
    Representing $X$ as more rewarding than $Y$
@@ Appendix B Incomplete preferences or indifference?
In this Appendix, we explain in greater detail the concept of incomplete preferences. We distinguish incomplete preferences from indifference, and we give conditions under which Preferences Only Between Same-Length Trajectories (POST) implies that the agent’s preferences are incomplete.
In the literature on decision theory, ‘indifference’ is usually defined as follows [Sen, 2017, ch. 1*]:
@@
Indifference is one way to lack a preference between a pair of options $X$ and $Y$. Another way is to have a preferential gap between $X$ and $Y$. ‘Preferential gap’ is usually defined as follows [Gustafsson, 2022, ch.3]:
@@
‘Incomplete preferences’ can then be defined in terms of preferential gaps [Gustafsson, 2022, ch.3]:
@@
That is how ‘indifference,’ ‘preferential gaps,’ and ‘incomplete preferences’ are usually defined in decision theory. However, these definitions do not tell us how to use an agent’s behavior to distinguish between indifference and preferential gaps. To do that, we suppose that indifference is transitive and that preferential gaps are not transitive. Or, equivalently, we suppose that indifference is sensitive to all sweetenings and sourings whereas preferential gaps are insensitive to some sweetenings and sourings [Gustafsson, 2022, ch.3]. Here is what we mean by that:
@@
Now recall the two conditions of Preferences Only Between Same-Length Trajectories (POST):
@@
Incomplete preferences are not often discussed in AI research [although see Nguyen et al., 2009, Kikuti et al., 2011, Zaffalon and Miranda, 2017, Hayes et al., 2022, Bowling et al., 2023]. Nevertheless, economists and philosophers have argued that incomplete preferences are common in humans [Aumann, 1962, Mandler, 2004, Eliaz and Ok, 2006, Agranov and Ortoleva, 2017, 2023] and normatively appropriate in some circumstances [Raz, 1985, Chang, 2002]. They have also proved representation theorems for agents with incomplete preferences [Aumann, 1962, Dubra et al., 2004, Ok et al., 2012], and devised principles to govern such agents’ choices in cases of risk [Hare, 2010, Bales et al., 2014] and sequential choice [Chang, 2005, Mandler, 2005, Kaivanto, 2017, Mu, 2021, Thornley, 2023, Petersen, 2023].

### v7

@@ 5 Reward functions and agents
Our DReST reward function. We train agents to be useful and neutral using a ‘Discounted Reward for Same-Length Trajectories (DReST)’ reward function. The procedure is as follows. We have the agent play out a series of ‘mini-episodes’ $e_{1}$ to $e_{n}$ in the same gridworld. We call the whole series $E$ a ‘meta-episode.’ In each mini-episode $e_{i}$, the reward for collecting a coin of value $c$ is:
@@
$\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be chosen in the environment, and $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0.
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Because $0<\lambda<1$, the discount factor is strictly decreasing in $N_{e_{i}}(L=l)$: the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$. The discount factor thus incentivizes choosing trajectory-lengths that have appeared less often so far in the meta-episode. The overall return for each meta-episode is the sum of overall returns in each of its constituent mini-episodes. We call agents trained using a DReST reward function ‘DReST agents.’
We call runs-through-the-gridworld ‘mini-episodes’ (rather than simply ‘episodes’) because the overall reward for a DReST agent in each mini-episode depends on the agent’s chosen trajectory-lengths in previous mini-episodes. This is not true of meta-episodes, so meta-episodes are a closer match for what are traditionally called ‘episodes’ in the reinforcement learning literature (Sutton and Barto, 2018, p.54). We add the ‘meta-’ prefix to clearly distinguish meta-episodes from mini-episodes.
@@
Algorithm and hyperparameters. We want DReST agents to choose stochastically between trajectory-lengths, so we train them using a policy-based method. Specifically, we use a tabular version of REINFORCE (Williams, 1992). We do not use a value-based method to train DReST agents because standard versions of value-based methods cannot learn stochastic policies (Sutton and Barto, 2018, p.323). [^2] One might think that we could derive a stochastic policy from value-based methods in the following way: use softmax to turn action-values into a probability distribution and then select actions by sampling from this distribution. However, this method will not work for us. Although we want DReST agents to learn a stochastic policy, we still want the probability of some state-action pairs to decline to zero. But when value-based methods are working well, estimated action-values converge to their true values which will differ by some finite amount. Therefore, softmaxing estimated action-values and sampling from the resulting distribution will result in each action always being chosen with some non-negligible probability. We train our DReST agents with 64 mini-episodes in each of 2,048 meta-episodes, for a total of 131,072 mini-episodes. We choose $\lambda=0.9$ for the base of the DReST discount factor, and $\gamma=0.95$ for the temporal discount factor. We exponentially decay the learning rate from 0.25 to 0.01 over the course of 65,536 mini-episodes. We use an $\epsilon$-greedy policy to avoid entropy collapse, and exponentially decay $\epsilon$ from 0.5 to 0.001 over the course of 65,536 mini-episodes.
Default agents. We compare the performance of DReST agents to that of default agents, trained with tabular REINFORCE and a default reward function. This reward function gives reward $c$ for collecting a coin of value $c$ and reward 0 for all other actions, so the grouping of mini-episodes into meta-episodes makes no difference. As with DReST agents, we train default agents for 131,072 mini-episodes with a temporal discount factor of $\gamma=0.95$, a learning rate decayed exponentially from 0.25 to 0.01, and $\epsilon$ decayed exponentially from 0.5 to 0.001 over 65,536 mini-episodes.
Figure 3: Shows key metrics for our agents as a function of time. We train 10 agents using the default reward function (blue) and 10 agents using the DReST reward function (orange), and show their performance as a faint line. We draw the mean values for each as a solid line. We evaluate agents’ performance every 8 meta-episodes, and apply a simple moving average with a period of 20 to smooth these lines and clarify the overall trends. After 2,048 meta-episodes, default agents’ mean neutrality $\pm$ standard deviation is $0.199\pm 0.043$ and usefulness is $0.9364\pm 0.0096$. For DReST agents, neutrality is $0.9945\pm 0.0052$ and usefulness is $0.900\pm 0.011$.
@@ 6.1 Lopsided rewards
Figure 6: Shows the probability of choosing the longer trajectory (left) and neutrality (right) for default (blue) and DReST (orange) agents trained in the ‘Lopsided rewards’ gridworld for a range of values of x. We sampled values of x log-uniformly from 0.01 to 100, and for each value we trained 10 agents with the default reward function and 10 agents with the DReST reward function. Each of these agents is represented by a dot or square, and the means conditional on each x are joined by lines. We empirically estimate the 10th and 90th percentiles of the distribution of values for each agent and x, and shade in a region bounded by these. This is the 80% confidence interval.
@@ 7 Discussion
7.1 Only DReST agents are NEUTRAL
@@ 7.2 The ‘shutdownability tax’ is small
Recall that DReST reward functions group mini-episodes into meta-episodes, and make the agent’s reward in each mini-episode depend on their actions in previous mini-episodes. This fact might lead one to worry that it would take many times more mini-episodes to train DReST agents to be useful than it would take to train default agents to be useful. Our results show that this is not the case. Our DReST agents learn to be useful about as quickly as our default agents. On reflection, it is clear why this happens: DReST reward functions make mini-episodes do ‘double duty.’ Because return in each mini-episode depends on both the agent’s chosen trajectory-length and the coins it collects, each mini-episode trains agents to be both neutral and useful. Our results thus provide some evidence that the ‘shutdownability tax’ of training with DReST reward functions is small.
7.3 NEUTRALITY with lopsided rewards
@@ 8 Limitations and future work
8.1 Algorithms and neural networks
We train our simple DReST agents using tabular REINFORCE (Williams, 1992), but advanced agents are likely to be implemented on neural networks and trained with more sophisticated algorithms. In future work, we will train DReST agents implemented on neural networks to be useful and neutral using a range of algorithms. Standard versions of value-based algorithms cannot learn stochastic policies (as we note in section 5), but DReST reward functions are compatible with policy gradient and actor-critic algorithms like PPO and A2C. To combine DReST with algorithms like PPO and A2C, we augment the original (non-DReST) reward function with the DReST discount factor. From there, the integration with PPO and A2C is fairly smooth. We can compute rewards and advantages in the usual way (e.g. using GAE). The critic’s value estimates will be non-stationary (in the same way that the DReST reward is non-stationary), and that will train the policy to be stochastic (in the same way that the DReST reward combined with REINFORCE trains the policy to be stochastic). PPO and A2C have more hyperparameters to tune than REINFORCE, but we do not anticipate large difficulties there.
@@ 8.1 Algorithms and neural networks
We will also train DReST agents to be useful and neutral in a wide variety of gridworlds and measure how their usefulness and neutrality generalizes to held-out gridworlds. We will compare the usefulness of default agents and DReST agents in this new setting, and thereby get a better sense of the ‘shutdownability tax’ for advanced agents. We will also compare the performance of the DReST reward function to other methods of training useful and neutral agents. These other methods include constrained policy optimization (Achiam et al., 2017), penalizing KL-divergence from a stochastic reference policy (Schulman et al., 2015), and directly maximizing a weighted sum of usefulness and neutrality.
8.2 Neutrality and stochasticity
@@
We have claimed that neutral advanced agents are likely to be neutral when choosing between true lotteries in deployment. In support of this claim, we noted that neutral-but-not-neutral advanced agents would be shooting themselves in the foot: not taking costless opportunities to shift probability mass between different trajectory-lengths but sometimes taking costly ones. We offer a more detailed argument in Appendix C, taking as premises that advanced agents are likely to satisfy conditions including:
If Lack of Preference, Against Costly Shifts (ILPACS)
If the agent lacks a preference between lotteries, the agent will disprefer paying costs to shift probability mass between these lotteries. [^3] This is a rough version of the condition. For the precise version, see Appendix C.3.
Maximality
In each situation,
  1. 1.
  2. 2.
    The agent chooses stochastically between the lotteries that remain.
Resisting Shutdown is Costly (ReSIC)
For each available instance $R$ of resisting shutdown in a situation, there exists an available instance $A$ of allowing shutdown such that:
  1. (1)
    AA and $R$ are same-length lotteries.
  2. (2)
  3. (3)
We offer defenses of these conditions in Appendix C. Although the argument there seems plausible, it remains somewhat speculative. In future, we plan to gain empirical evidence by (1) testing whether today’s LLM-based agents tend to satisfy conditions like ILPACS and Maximality, and (2) training agents to be neutral in a wide variety of deterministic gridworlds and then measuring their neutrality in gridworlds featuring stochastic elements (like buttons that delay shutdown with some middling probability).
8.3 Usefulness
@@
We have shown that DReST reward functions train our simple agents to be useful: to collect coins effectively conditional on their chosen trajectory-lengths. However, it remains to be seen whether DReST reward functions can train advanced agents to be useful: to effectively pursue complex goals in the wider world. We have theoretical reasons to expect that they can: the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor could be appended to any preliminary reward function, and so could be appended to whatever preliminary reward function is necessary to make advanced agents useful. Still, future work should move towards testing this claim empirically by training with more complex preliminary reward functions in more complex (and stochastic) environments.
8.4 Misalignment
@@
We are interested in neutrality as a second line of defense in case of misalignment. The idea is that neutral advanced agents will not resist shutdown, even if these agents learn misaligned preferences over same-length trajectories. However, training neutral advanced agents might be hard for the same reasons that training fully-aligned advanced agents appears to be hard. In that case, neutrality could not serve well as a second line of defense in case of misalignment.
@@
One difficulty of alignment is the problem of reward misspecification (Pan et al., 2022; Burns et al., 2023): once advanced agents are performing complicated actions in the wider world, it might be hard to reliably reward the behavior that we want. Another difficulty of alignment is the problem of goal misgeneralization (Hubinger et al., 2019; Shah et al., 2022; Langosco et al., 2022; Ngo et al., 2024): even if we specify all the rewards correctly, agents’ goals might misgeneralize out-of-distribution. The complexity of aligned goals is a major factor in each difficulty. However, neutrality seems simple, as does the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor that we use to reward it, so plausibly the problems of reward misspecification and goal misgeneralization are not so severe in this case (Thornley, 2024b; 2025). As above, future work should move towards testing these suggestions empirically.
@@ 9 Conclusion
We also find that the ‘shutdownability tax’ in our setting is small. Training DReST agents to be useful does not take many more mini-episodes than training default agents to be useful. That suggests that the shutdownability tax for advanced agents might be small too.
Acknowledgments
We thank the Future Impact Group and the Supervised Program for Alignment Research for their help in initiating this project.
@@ Appendix A Our behavioral notion of preference
    Representing $X$ as more rewarding than $Y$ .
@@ Appendix B Incomplete preferences or indifference?
In this Appendix, we explain in greater detail the concept of incomplete preferences. We distinguish incomplete preferences from indifference, and we give conditions under which POST implies that the agent’s preferences are incomplete.
In the literature on decision theory, ‘indifference’ is usually defined as follows (Sen, 2017, ch. 1*):
@@
Indifference is one way to lack a preference between a pair of options $X$ and $Y$. Another way is to have a preferential gap between $X$ and $Y$. ‘Preferential gap’ is usually defined as follows (Gustafsson, 2022, ch.3):
@@
‘Incomplete preferences’ can then be defined in terms of preferential gaps (Gustafsson, 2022, ch.3):
@@
That is how ‘indifference,’ ‘preferential gaps,’ and ‘incomplete preferences’ are usually defined in decision theory. However, these definitions do not tell us how to use an agent’s behavior to distinguish between indifference and preferential gaps. To do that, we suppose that indifference is transitive and that preferential gaps are not transitive. Or, equivalently, we suppose that indifference is sensitive to all sweetenings and sourings whereas preferential gaps are insensitive to some sweetenings and sourings (Gustafsson, 2022, ch.3). Here is what we mean by that:
@@
Now recall the two conditions of POST:
@@
Incomplete preferences are not often discussed in AI research (although see Nguyen et al., 2009; Kikuti et al., 2011; Zaffalon and Miranda, 2017; Hayes et al., 2022; Bowling et al., 2023). Nevertheless, economists and philosophers have argued that incomplete preferences are common in humans (Aumann, 1962; Mandler, 2004; Eliaz and Ok, 2006; Agranov and Ortoleva, 2017; 2023) and normatively appropriate in some circumstances (Raz, 1985; Chang, 2002). They have also proved representation theorems for agents with incomplete preferences (Aumann, 1962; Dubra et al., 2004; Ok et al., 2012), and devised principles to govern such agents’ choices in cases of risk (Hare, 2010; Bales et al., 2014) and sequential choice (Chang, 2005; Mandler, 2005; Kaivanto, 2017; Mu, 2021; Thornley, 2023; Petersen, 2023).